# Reinforcement Learning from Probabilistic Forecasts for Safe Decision-Making via Conditional Value-at-Risk Planning

### v1

@@ Abstract
Sequential decisions in volatile, high-stakes settings require more than maximizing expected return; they require principled uncertainty management. This paper presents the Uncertainty-Aware Markov Decision Process (UAMDP), a unified framework that couples Bayesian forecasting, posterior-sampling reinforcement learning, and planning under a conditional value-at-risk (CVaR) constraint. In a closed loop, the agent updates its beliefs over latent dynamics, samples plausible futures via Thompson sampling, and optimizes policies subject to preset risk tolerances. We establish regret bounds that converge to the Bayes-optimal benchmark under standard regularity conditions. We evaluate UAMDP in two domains—high-frequency equity trading and retail inventory control—both marked by structural uncertainty and economic volatility. Relative to strong deep learning baselines, UAMDP improves long-horizon forecasting accuracy (RMSE decreases by up to 25% and sMAPE by 32%), and these gains translate into economic performance: the trading Sharpe ratio rises from 1.54 to 1.74 while maximum drawdown is roughly halved. These results show that integrating calibrated probabilistic modeling, exploration aligned with posterior uncertainty, and risk-aware control yields a robust, generalizable approach to safer and more profitable sequential decision-making.
@@ 1 Introduction
Robust decision-making under uncertainty is critical in dynamic, high-stakes domains such as finance, healthcare, and energy systems, where even marginally suboptimal actions can incur substantial economic or human costs. These environments are not only volatile but also structurally complex, often exhibiting latent dynamics that challenge both predictive accuracy and policy reliability. As such, effective decision frameworks must go beyond deterministic predictions and pointwise optimization to integrate mechanisms for uncertainty quantification and risk-sensitive control [1].
@@
Despite significant advances, most traditional reinforcement learning (RL) approaches to forecasting and decision-making remain focused on maximizing expected returns. This narrow objective often yields policies that are overconfident in their assumptions and brittle in distributional shifts or unanticipated conditions. Meanwhile, state-of-the-art forecasting models frequently produce deterministic or under-calibrated outputs, diminishing their value for downstream decisions that depend on uncertainty awareness [2]. These parallel deficiencies highlight a fundamental gap: the lack of a cohesive, uncertainty-aware loop that connects probabilistic modeling with sequential control.
@@
Recent progress in probabilistic forecasting, particularly using Gaussian processes (GPs) and Temporal Fusion Transformers (TFTs), has demonstrated the feasibility of generating calibrated predictive intervals, offering a more nuanced view of future states [2].
@@
Yet in practice, these models are rarely integrated into the control layer of forecasting software. Instead, they are treated as peripheral inputs, decoupled from the agent’s learning and planning mechanisms. In parallel, deep RL has demonstrated remarkable performance on complex tasks but often relies on heuristic exploration and disregards uncertainty estimates altogether [3]. Consequently, learned policies may perform well under familiar conditions but collapse when confronted with unfamiliar inputs, precisely where uncertainty-aware reasoning is most needed [4].
@@
Moreover, mainstream RL formulations frequently assume fully observable and discretized environments. This abstraction neglects the partial observability and continuous dynamics that characterize many real-world systems, from patient health trajectories to financial markets [5]. Even within methods that nominally address uncertainty, the treatment is often superficial, relying on simplistic approximations or failing to propagate uncertainty through the decision pipeline. This methodological gap severely limits the reliability of RL-based systems in environments where uncertainty is inherent and safety, robustness, and risk awareness are not optional but essential [6].
@@
To bridge these gaps, we propose the Uncertainty-Aware Markov Decision Process (UAMDP), a novel framework that unifies probabilistic forecasting and RL in a closed-loop, Bayes-adaptive MDP formulation. The UAMDP maintains a posterior belief over latent environment parameters, which is continuously updated using forecasts from state-of-the-art probabilistic models-GPs for moderate-scale problems and Monte Carlo dropout transformers for high-dimensional time series. These beliefs drive posterior-sampling RL, enabling the agent to simulate plausible futures and optimize policies that are both exploratory and risk-aware [7]. The primary contributions of this work are summarized below:
@@ 2.1 Probabilistic time-series forecasting
Bayesian approaches have long emphasized the role of predictive distributions in sequential decision-making. GP models have advanced non-parametric forecasting by offering analytically tractable uncertainty quantification, establishing themselves as foundational tools in high-volatility domains such as finance and critical care [8].
The emergence of deep learning has significantly extended the capabilities of probabilistic forecasting, enabling accurate modeling across large-scale and multi-series datasets. Frameworks such as DeepAR, which employ autoregressive recurrent neural networks (RNNs) to directly estimate likelihood parameters, have achieved notable accuracy gains, reportedly reducing forecast error by up to 15% in domains such as retail and energy [9, 10, 11]. Further advances, including TFTs with Monte Carlo dropout [2] and multi-horizon quantile RNNs [12], have enriched forecast calibration, while recent regression-based models address interval misestimation through post-hoc correction.
Latent-state forecasting frameworks have also shown considerable promise by integrating deep neural architectures with Bayesian filtering. Variational RNNs [13], deep Markov models [14], and deep state-space models [15] capture latent regime transitions that often evade purely deterministic models. Complementary approaches, such as Bayesian ensembling and model stacking, alongside classical methods such as quantile regression and interval prediction [16], further reduce model risk. Collectively, these contributions underscore the necessity of generating credible predictive intervals, not only to improve forecasting accuracy but to enable safe and informed downstream decision-making under uncertainty.
@@ 2.2 Reinforcement learning under uncertainty
RL research has long addressed the inherent challenges posed by stochastic rewards and model uncertainty. Probably Approximately Correct (PAC) MDP algorithms, such as R-MAX [17] and UCRL [18], offer finite-sample performance guarantees by leveraging optimism in the face of uncertainty. More recent advances in distributional RL model the full return distribution rather than expected values, enabling richer value representations [19]. In parallel, risk-sensitive RL has introduced tail-aware optimization objectives such as CVaR, replacing the conventional focus on expected returns with measures of downside exposure [20].
@@
Domain-specific studies have further demonstrated the practical importance of uncertainty quantification. In autonomous driving, posterior variance estimates are used to improve collision-avoidance strategies, particularly in high-risk scenarios such as intersection navigation [21]. In finance, uncertainty-aware policies have incorporated variational autoencoder-based epistemic penalties to enhance portfolio resilience, complemented by hierarchical controllers and normalizing-flow transition models capable of adapting to market shocks [22, 23]. Despite these advances, recent reviews underscore persistent barriers to industrial deployment—most notably, limited robustness under distributional shift and the absence of standardized evaluation protocols [24].
@@
Bayesian RL formalizes exploration–exploitation trade-offs via probabilistic inference [25]. Posterior-sampling RL, achieving near-optimal regret, has gained attention due to its effective policy optimization [7]. Extensions such as Bayesian Q-learning [26], variance-adaptive Thompson sampling [27], and meta-contextual bandits broaden applicability [28]. The probabilistic inference for learning control (PILCO) framework notably demonstrates data-efficient Bayesian policy learning by leveraging GP moment propagation [29], with Monte Carlo and dropout-based variants addressing partial observability and global stability.
@@
Recent contributions to Bayesian RL include Bayesian exploration networks (BEN), which enable model-free learning of value distributions under uncertainty, and Bayesian residual policy optimization, which incorporates expert priors to refine action selection [30]. These innovations represent substantial progress toward Bayes-optimal decision-making, particularly in high-stakes settings where uncertainty plays a central role. Notably, the integration of probabilistic forecasting with RL has demonstrated promising results across a range of application domains.
@@
In finance, DeepAR-based predictive distributions have been employed to guide portfolio optimization, yielding substantial improvements in risk-adjusted returns [31]. In healthcare, Bayesian methods have enhanced intensive care unit risk scoring and sepsis treatment through calibrated probabilistic forecasting [6] and have been further applied to personalized insulin dosing using CVaR-constrained optimization [32]. In autonomous systems, forecast-informed RL has improved decision robustness under sensor noise and uncertainty, reducing accident risk and enhancing navigational performance [21]. Similarly, energy management systems have benefited from multiscale forecasting integrated with RL, enabling more effective demand response strategies and reducing peak load stress [33].
@@ 2.3 Integrated forecasting and control
Classical model-predictive control (MPC) has long incorporated forecast information to guide horizon-based decision-making [34]. Contemporary extensions, including stochastic and chance-constrained MPC, explicitly model uncertainty to enhance control robustness [35]. Deep RL approaches have further advanced this integration. Methods such as probabilistic ensembles with trajectory sampling (PETS) combine probabilistic ensemble dynamics models with trajectory-sampling MPC, achieving competitive performance with substantially improved sample efficiency [36].
More recent frameworks, such as model predictive path integral control [37] and Dreamer-V2 [38], expand this paradigm by leveraging learned dynamics models for real-time control. Bayesian adaptive deep RL methods integrate neural architectures with Thompson sampling, providing strong empirical performance alongside theoretical regret guarantees. Collectively, these approaches underscore the value of explicitly modeling uncertainty to achieve efficient, high-performance, and robust decision making.
@@ 3 Uncertainty-aware Markov decision process
$S,A$ State and action spaces
@@
$\theta\in\Theta$ Latent environment/forecast parameters
@@
We write $\theta=(\phi,\psi)$, where $\phi$ indexes the latent transition kernel $P_{\phi}\!\left(s^{\prime}\mid s,a,x_{t}\right)$ and $\psi$ indexes the forecaster’s predictive mean/variance. Exogenous covariates $x_{t}$ are observable and thus excluded from $\theta$.
@@ 3.1 Procedure flow
Let $p_{\phi}(x_{t+1}\mid s_{t},a_{t},\theta)$ denote a probabilistic one-step forecaster. For moderate feature dimension $d$, we instantiate $p_{\phi}$ as a GP; when $d$ is large, we employ a TFT equipped with Monte-Carlo dropout so that every predictive call returns a full distribution rather than a point estimate. The forecaster thus supplies the likelihood term required for Bayesian inference over the latent environment parameter $\theta$.
@@
Initially, the procedure instantiates its three core components. It sets a prior belief $b_{0}(\theta)$ over the latent environment parameter; selects the forecaster $p_{\phi}(x_{t+1}\mid s_{t},a_{t},\theta)$, implemented as a GP when the feature dimension $d$ is moderate or as a TFT with Monte-Carlo dropout when $d$ is high; and loads the policy/critic weights $\psi$. These elements jointly define the agent’s initial information state.
@@
Next, at the start of each episode, the agent performs Thompson exploration by sampling a concrete model $\theta_{k}\sim b_{0}(\theta)$. This random draw aligns the forthcoming policy with the agent’s current epistemic uncertainty, thereby encouraging information-seeking behavior when the prior is diffuse. Planning then unfolds over a finite horizon $H$. Conditioned on the sampled $\theta_{k}$ and the current hyper-state $(b_{t},s_{t})$, the agent applies model-predictive control, Monte Carlo tree search, or differentiable rollouts to maximize the Bayesian action value:
$Q(b_{t},s_{t},a)=\mathbb{E}_{\theta\sim b_{t}}\!\left[r(s_{t},a)+\gamma\,V^{\pi}\!\big(b_{t+1}^{(\theta)},s_{t+1}\big)\right].$
such that parameter values explaining the data gain posterior mass, and future decisions become correspondingly sharper. Optionally, a risk-sensitive layer is interposed before action execution. In that case, the planner replaces the mean-based utility with a coherent risk measure, e.g., evaluates $\mathrm{CVaR}_{\alpha}\!\left[Z_{\psi}(b_{t},s_{t},a)\right]$ in lieu of $Q$, or enforces chance constraints:
thereby guaranteeing tail-risk mitigation or probabilistic safety. Finally, the loop of sampling, planning, execution, and Bayesian updating repeats until the fixed horizon $T$ is reached or an earlier task-specific termination criterion is met. Through this closed-loop interaction, the UAMDP continuously couples uncertainty-aware forecasting with decision-making, ensuring that every action both exploits current knowledge and refines future understanding.
@@
2 Load forecaster $(F)$ weights $\phi$ and policy weights $\psi$;
3 for each episode $k$ do

### v2

# UAMDP: Uncertainty-Aware Markov Decision Process for Risk-Constrained Reinforcement Learning from Probabilistic Forecasts

@@ Abstract
Sequential decisions in volatile, high-stakes settings require more than maximizing expected return; they require principled uncertainty management. This paper presents the Uncertainty-Aware Markov Decision Process (UAMDP), a unified framework that couples Bayesian forecasting, posterior-sampling reinforcement learning, and planning under a conditional value-at-risk (CVaR) constraint. In a closed loop, the agent updates its beliefs over latent dynamics, samples plausible futures via Thompson sampling, and optimizes policies subject to preset risk tolerances. We establish regret bounds that converge to the Bayes-optimal benchmark under standard regularity conditions. We evaluate UAMDP in two domains including high-frequency equity trading and retail inventory control, both marked by structural uncertainty and economic volatility. Relative to strong deep learning baselines, UAMDP improves long-horizon forecasting accuracy (RMSE decreases by up to 25% and sMAPE by 32%), and these gains translate into economic performance: the trading Sharpe ratio rises from 1.54 to 1.74 while maximum drawdown is roughly halved. These results show that integrating calibrated probabilistic modeling, exploration aligned with posterior uncertainty, and risk-aware control yields a robust, generalizable approach to safer and more profitable sequential decision-making.
@@ 1 Introduction
Robust decision-making under uncertainty is critical in dynamic, high-stakes domains such as finance, healthcare, and energy systems, where even marginally suboptimal actions can incur substantial economic or human costs. These environments are not only volatile but also structurally complex, often exhibiting latent dynamics that challenge both predictive accuracy and policy reliability. As such, effective decision frameworks must go beyond deterministic predictions and pointwise optimization to integrate mechanisms for uncertainty quantification and risk-sensitive control [hsieh2025risk].
@@
Despite significant advances, most traditional reinforcement learning (RL) approaches to forecasting and decision-making remain focused on maximizing expected returns. This narrow objective often yields policies that are overconfident in their assumptions and brittle in distributional shifts or unanticipated conditions. Meanwhile, state-of-the-art forecasting models frequently produce deterministic or under-calibrated outputs, diminishing their value for downstream decisions that depend on uncertainty awareness [lim2021temporal]. These parallel deficiencies highlight a fundamental gap: the lack of a cohesive, uncertainty-aware loop that connects probabilistic modeling with sequential control.
@@
Recent progress in probabilistic forecasting, particularly using Gaussian processes (GPs) and Temporal Fusion Transformers (TFTs), has demonstrated the feasibility of generating calibrated predictive intervals, offering a more nuanced view of future states [lim2021temporal].
@@
Yet in practice, these models are rarely integrated into the control layer of forecasting software. Instead, they are treated as peripheral inputs, decoupled from the agent’s learning and planning mechanisms. In parallel, deep RL has demonstrated remarkable performance on complex tasks but often relies on heuristic exploration and disregards uncertainty estimates altogether [sutton1998reinforcement]. Consequently, learned policies may perform well under familiar conditions but collapse when confronted with unfamiliar inputs, precisely where uncertainty-aware reasoning is most needed [abdar2021review].
@@
Moreover, mainstream RL formulations frequently assume fully observable and discretized environments. This abstraction neglects the partial observability and continuous dynamics that characterize many real-world systems, from patient health trajectories to financial markets [ladosz2022exploration]. Even within methods that nominally address uncertainty, the treatment is often superficial, relying on simplistic approximations or failing to propagate uncertainty through the decision pipeline. This methodological gap severely limits the reliability of RL-based systems in environments where uncertainty is inherent and safety, robustness, and risk awareness are not optional but essential [komorowski2018artificial].
@@
To bridge these gaps, we propose the Uncertainty-Aware Markov Decision Process (UAMDP), a novel framework that unifies probabilistic forecasting and RL in a closed-loop, Bayes-adaptive MDP formulation. The UAMDP maintains a posterior belief over latent environment parameters, which is continuously updated using forecasts from state-of-the-art probabilistic models-GPs for moderate-scale problems and Monte Carlo dropout transformers for high-dimensional time series. These beliefs drive posterior-sampling RL, enabling the agent to simulate plausible futures and optimize policies that are both exploratory and risk-aware [osband2013more]. The primary contributions of this work are summarized below:
@@ 2.1 Probabilistic time-series forecasting
Bayesian approaches have long emphasized the role of predictive distributions in sequential decision-making. GP models have advanced non-parametric forecasting by offering analytically tractable uncertainty quantification, establishing themselves as foundational tools in high-volatility domains such as finance and critical care [liu2021incremental]. Recent work on dynamic feature-fusion architectures further extends these ideas to large-scale time series with rich covariates, for example the DFF-Net model of [xiao2025dff], which learns to adaptively combine multiple feature channels over time.
The emergence of deep learning has significantly extended the capabilities of probabilistic forecasting, enabling accurate modeling across large-scale and multi-series datasets. Frameworks such as DeepAR, which employ autoregressive recurrent neural networks (RNNs) to directly estimate likelihood parameters, have achieved notable accuracy gains, reportedly reducing forecast error by up to 15% in domains such as retail and energy [salinas2020deepar, koren2022inventory, koren2024inventory]. Further advances, including TFTs with Monte Carlo dropout [lim2021temporal] and multi-horizon quantile RNNs [gasthaus2019probabilistic], have enriched forecast calibration, while recent regression-based models address interval misestimation through post-hoc correction. Hybrid approaches that combine data-driven models with structured prior knowledge offer another promising direction for robust probabilistic forecasting [chen2024towards].
Latent-state forecasting frameworks have also shown considerable promise by integrating deep neural architectures with Bayesian filtering. Variational RNNs [chung2015gated], deep Markov models [fraccaro2016sequential], and deep state-space models [rangapuram2018deep] capture latent regime transitions that often evade purely deterministic models. Complementary approaches, such as Bayesian ensembling and model stacking, alongside classical methods such as quantile regression and interval prediction [kabir2018partial], further reduce model risk. Related work in approximate reasoning has studied imprecise and non-stationary uncertainty models for sequential predictions [persiau2022dis], emphasizing the importance of credible interval estimates. Collectively, these contributions underscore the necessity of generating credible predictive intervals, not only to improve forecasting accuracy but to enable safe and informed downstream decision-making under uncertainty.
@@ 2.2 Reinforcement learning under uncertainty
RL research has long addressed the inherent challenges posed by stochastic rewards and model uncertainty. Probably Approximately Correct (PAC) MDP algorithms, such as R-MAX [brafman2002r] and UCRL [auer2008near], offer finite-sample performance guarantees by leveraging optimism in the face of uncertainty. More recent advances in distributional RL model the full return distribution rather than expected values, enabling richer value representations [dabney2018distributional]. In parallel, risk-sensitive RL has introduced tail-aware optimization objectives such as CVaR, replacing the conventional focus on expected returns with measures of downside exposure [tamar2015optimizing]. Complementary lines of work in approximate reasoning and decision theory study risk-sensitive Markov decision processes with budget constraints [moreira2024efficient] and three-way decision models that explicitly encode different attitudes toward risk and uncertainty [xin2025interval, li2025improved, suo2024review].
@@
Domain-specific studies have further demonstrated the practical importance of uncertainty quantification. In autonomous driving, posterior variance estimates are used to improve collision-avoidance strategies, particularly in high-risk scenarios such as intersection navigation [hoel2020reinforcement]. In finance, uncertainty-aware policies have incorporated variational autoencoder-based epistemic penalties to enhance portfolio resilience, complemented by hierarchical controllers and normalizing-flow transition models capable of adapting to market shocks [enkhsaikhan2024uncertainty, koren2023forecasting]. Despite these advances, recent reviews underscore persistent barriers to industrial deployment—most notably, limited robustness under distributional shift and the absence of standardized evaluation protocols [hambly2023recent].
@@
Bayesian RL formalizes exploration–exploitation trade-offs via probabilistic inference [ghavamzadeh2015bayesian]. Posterior-sampling RL, achieving near-optimal regret, has gained attention due to its effective policy optimization [osband2013more]. Extensions such as Bayesian Q-learning [dearden1998bayesian], variance-adaptive Thompson sampling [thompson1933likelihood], and meta-contextual bandits broaden applicability [koren2024automated]. The probabilistic inference for learning control (PILCO) framework notably demonstrates data-efficient Bayesian policy learning by leveraging GP moment propagation [deisenroth2011pilco], with Monte Carlo and dropout-based variants addressing partial observability and global stability.
@@
Recent contributions to Bayesian RL include Bayesian exploration networks (BEN), which enable model-free learning of value distributions under uncertainty, and Bayesian residual policy optimization, which incorporates expert priors to refine action selection [lee2021bayesian]. These innovations represent substantial progress toward Bayes-optimal decision-making, particularly in high-stakes settings where uncertainty plays a central role. In parallel, learning human-interpretable decision catalogues and scoring systems for situated decision making has attracted renewed interest [heid2024learning], reflecting a broader shift toward transparent policies. Notably, the integration of probabilistic forecasting with RL has demonstrated promising results across a range of application domains.
@@
In finance, DeepAR-based predictive distributions have been employed to guide portfolio optimization, yielding substantial improvements in risk-adjusted returns [wen2017multi]. In healthcare, Bayesian methods have enhanced intensive care unit risk scoring and sepsis treatment through calibrated probabilistic forecasting [komorowski2018artificial] and have been further applied to personalized insulin dosing using CVaR-constrained optimization [lu2023structured]. In autonomous systems, forecast-informed RL has improved decision robustness under sensor noise and uncertainty, reducing accident risk and enhancing navigational performance [hoel2020reinforcement]. Similarly, energy management systems have benefited from multiscale forecasting integrated with RL, enabling more effective demand response strategies and reducing peak load stress [yang2025derl].
@@ 2.3 Integrated forecasting and control
Classical model-predictive control (MPC) has long incorporated forecast information to guide horizon-based decision-making [rawlings2020model]. Contemporary extensions, including stochastic and chance-constrained MPC, explicitly model uncertainty to enhance control robustness [kouvaritakis2015model]. Deep RL approaches have further advanced this integration. Methods such as probabilistic ensembles with trajectory sampling (PETS) combine probabilistic ensemble dynamics models with trajectory-sampling MPC, achieving competitive performance with substantially improved sample efficiency [chua2018deep].
More recent frameworks, such as model predictive path integral control [williams2017model] and Dreamer-V2 [hafner2019dream], expand this paradigm by leveraging learned dynamics models for real-time control. Bayesian adaptive deep RL methods integrate neural architectures with Thompson sampling, providing strong empirical performance alongside theoretical regret guarantees. These developments align with broader efforts in approximate reasoning and decision theory to design robust decision rules under uncertainty [suo2024review, heid2024learning]. Collectively, these approaches underscore the value of explicitly modeling uncertainty to achieve efficient, high-performance, and robust decision making.
@@ 3 Uncertainty-aware Markov decision process
$\mathcal{S},\mathcal{A}$ State and action spaces
@@
$\theta\in\Theta$ Latent environment parameters
@@
We write $\theta$ to index latent environment parameters governing the transition kernel $P_{\theta}(s^{\prime}\mid s,a)$. The probabilistic forecaster is parameterized by weights $\phi$, and the policy and distributional critic are parameterized by weights $\psi$; during online interaction, $\phi$ and $\psi$ are treated as fixed and the only online model update acts on the belief $b_{t}(\theta)$. Since the augmented state $s_{t}=(x_{1:t},a_{0:t-1})$ contains the observable history, all dependence on the current observation is captured through $s_{t}$. Concretely, in UAMDP the kernel $P_{\theta}$ is induced by the forecaster: given $(s_{t},a_{t},\theta)$, we draw $x_{t+1}\sim p_{\phi}(\cdot\mid s_{t},a_{t},\theta)$ and set $s_{t+1}=\tau(s_{t},a_{t},x_{t+1}):=(x_{1:t+1},a_{0:t})$.
@@ 3.1 Procedure flow
Let $p_{\phi}(x_{t+1}\mid s_{t},a_{t},\theta)$ denote a probabilistic one-step forecaster for the next observation $x_{t+1}\in\mathbb{R}^{d}$. For moderate observation dimension $d$, we instantiate $p_{\phi}$ as a GP; when $d$ is large, we employ a TFT equipped with Monte-Carlo dropout so that every predictive call returns a full predictive distribution rather than a point estimate. The forecaster thus supplies the likelihood term required for Bayesian inference over the latent environment parameter $\theta$.
Specifically, for moderate observation dimension $d$, we instantiate $p_{\phi}(x_{t+1}\mid s_{t},a_{t},\theta)$ via GP regression as follows. Recall that $\theta\in\Theta$ denotes latent environment parameters (fixed within an episode) that govern the induced transition kernel $P_{\theta}(s^{\prime}\mid s,a)$ defined above, while $\phi$ denotes forecaster parameters obtained by offline pre-training and then held fixed during RL training. To define the GP, let $u_{t}:=f(s_{t},a_{t})\in\mathbb{R}^{p}$ be a fixed feature map of the observable history and action, and form an augmented GP input $z_{t}:=\mathrm{concat}(u_{t},\eta(\theta))\in\mathbb{R}^{p+p_{\theta}}$, where $\eta(\theta)$ is an encoding of $\theta$ (e.g., identity for continuous $\theta$, or one-hot/embedding for discrete $\theta$). For each output coordinate $j\in\{1,\dots,d\}$, we place a GP prior on a latent function $g_{j}(\cdot)\sim\mathcal{GP}\!\big(m_{\phi,j}(\cdot),k_{\phi,j}(\cdot,\cdot)\big)$ and assume the observation model
$x_{t+1}^{(j)}=g_{j}(z_{t})+\varepsilon_{t}^{(j)},\qquad\varepsilon_{t}^{(j)}\sim\mathcal{N}(0,\sigma_{\phi,j}^{2}).$
After offline training (absorbed into $\phi$), the resulting one-step predictive distribution used in UAMDP is
$p_{\phi}(x_{t+1}\mid s_{t},a_{t},\theta)=\prod_{j=1}^{d}\mathcal{N}\!\big(x_{t+1}^{(j)};\mu_{\phi,j}(z_{t}),\,\Sigma_{\phi,j}(z_{t})\big),$
where $\mu_{\phi,j}(z_{t})$ and $\Sigma_{\phi,j}(z_{t})$ are the standard GP posterior mean and variance (with fixed hyperparameters absorbed into $\phi$); the product form corresponds to independent-output GPs for tractability. This density provides the likelihood term in the Bayes update of $b_{t}(\theta)$. Here $\phi$ absorbs the GP specification learned offline, including the mean and kernel hyperparameters (e.g., length-scales and output variance) and the observation-noise variance $\sigma_{\phi,j}^{2}$, as well as any fixed sparsification choices (e.g., inducing inputs, if used). The functions $\mu_{\phi,j}(z_{t})$ and $\Sigma_{\phi,j}(z_{t})$ are obtained from the standard GP posterior predictive equations conditioned on the offline training data (and, if applicable, the inducing representation), so that $p_{\phi}(\cdot\mid s_{t},a_{t},\theta)$ is fully specified at test time and the forecaster parameters $\phi$ are not updated online. Given the history-state definition $s_{t}=(x_{1:t},a_{0:t-1})$, the induced state-transition kernel is obtained by first sampling $x_{t+1}\sim p_{\phi}(\cdot\mid s_{t},a_{t},\theta)$ and then setting $s_{t+1}=\tau(s_{t},a_{t},x_{t+1})=(x_{1:t+1},a_{0:t})$, which defines $P_{\theta}(s_{t+1}\mid s_{t},a_{t})$.
In practice, this choice reflects a trade-off between calibration and computational cost rather than a hard requirement of UAMDP. Following the common use of Gaussian processes in moderate-scale settings, we instantiate $p_{\phi}$ as a GP only when both the observation dimension and the available sample size are modest, so that training and prediction remain tractable. For high-dimensional or large multi-series problems, we instead employ TFT with Monte-Carlo dropout, as discussed in the empirical section. In that regime, a GP becomes less efficient and more sensitive to irrelevant covariates, whereas TFT can learn latent representations and capture long-range dependencies across many input channels. Monte-Carlo dropout at inference time then induces a predictive distribution over $x_{t+1}$ rather than a single point forecast, and this distribution is plugged into the Bayesian update as the likelihood term. Importantly, the surrounding UAMDP loop including belief updating, Thompson sampling, and planning remains unchanged; only the underlying probabilistic forecaster $p_{\phi}$ is swapped.
@@
Initially, the procedure instantiates its three core components. It sets a prior belief $b_{0}(\theta)$ over the latent environment parameter; selects the forecaster $p_{\phi}(x_{t+1}\mid s_{t},a_{t},\theta)$, implemented as a GP when the observation dimension $d$ is moderate or as a TFT with Monte-Carlo dropout when $d$ is high; and loads the policy/critic weights $\psi$. These elements jointly define the agent’s initial information state.
@@
Next, at the start of each episode, the agent performs Thompson exploration by sampling a concrete model $\theta_{k}\sim b(\theta)$, where $b(\theta)$ denotes the current belief (initialized as $b_{0}(\theta)$). This random draw aligns the forthcoming policy with the agent’s current epistemic uncertainty, thereby encouraging information-seeking behavior when the belief is diffuse. At each decision step, planning uses a finite lookahead horizon $H$. Conditioned on the sampled $\theta_{k}$ and the current hyper-state $(b_{t},s_{t})$, the agent applies a planner to maximize the Bayesian action value; in implementation, $\theta_{k}$ provides a Monte-Carlo approximation for evaluating the belief expectation inside the planner (e.g., via rollouts under $P_{\theta_{k}}$). In our empirical study (Section 4), this planner is instantiated as a bounded-depth Monte Carlo tree search (MCTS); the model-predictive control (MPC) and differentiable-rollout backends shown in Fig. 1 illustrate that the UAMDP framework is compatible with alternative trajectory-optimization controllers, but they are not used for the experiments reported here. The Bayesian action value is given by
$Q(b_{t},s_{t},a)=\mathbb{E}_{\theta\sim b_{t}}\!\left[r(s_{t},a)+\gamma\,V^{\pi}\!\big(b_{t+1}^{(\theta)},s_{t+1}\big)\right],$
where $V^{\pi}$ denotes the value induced by the rollout/evaluation policy used by the planner (e.g., the default policy in MCTS).
Figure 1: Flow of the proposed framework.
@@
2 Load forecaster $(F)$ weights $\phi$ and policy/critic weights $\psi$;
$n\leftarrow 0$;
// global interaction step
3 while $n<T$ do